Waterbury Insurance Company wants to study the relationship between the amount of fire damage and the distance between the bub. How much damage would you estimate for a fire 4 miles from the nearest fire station? (Round your answer to the nearest dolc-1. Determine and interpret the coefficient of determination. (Round your answer to 3 decimal places.) Coefficient of determc-2. Fill in the blank below. (Round your answer to one decimal place.) % of the variation in damage is explained by variatiod-1. Determine the correlation coefficient. (Round your answer to 3 decimal places.) Correlation coefficiente-2. Compute the value of the test statistic. (Round your answer to 2 decimal places.) Value of the test statistic
Answer #1

a-1. y = 13.760 +3.771 be for so Estimated Damage = 13.76 + 3•7184 X= 4 Estimated CD Coefficient of Damage 28.844 Determinati